OpenStudy (anonymous):

For a party, you are cooking a large amount of stew that has meat, potatoes, and carrots. The meat costs $6 per pound, the potatoes cost $3 per pound, and the carrots cost $1 per pound. You spend $48.50 on 13.5 pounds of food. You buy twice as many carrots as potatoes. ---------------------------------- I need 3 equations that represent how much food I bought (in a triple variable systematic way)

OpenStudy (anonymous):

Let x be number of pounds of meat

OpenStudy (anonymous):

let x = lb of meat y = lb of potatoes z = lb of carrots x + y + z = 13.5 6x + 3y + z = 48.5 z = 2y x + y + 2y = 13.5 x + 3y = 13.5 x = 13.5-3y 6(13.5-3y) + 3y + 2y = 48.5 81 - 18y + 5y = 48.5 81 - 48.5 = 13y 32.5 = 13y 2.5 = y x = 13.5 - 3y x = 13.5 -3*2.5 x = 13.5 - 7.5 x = 6 z = 2y z = 2*2.5 z = 5